# From square plaquettes to triamond lattices for SU(2) gauge theory

### v1

@@ Abstract
Lattice gauge theory should be able to address significant new scientific questions when implemented on quantum computers. In practice, error-mitigation techniques have already allowed encouraging progress on small lattices. In this work we focus on a truncated version of SU(2) gauge theory, which is a familiar non-Abelian step toward quantum chromodynamics. First, we demonstrate effective error mitigation for imaginary time evolution on a lattice having two square plaquettes, obtaining the ground state using an IBM quantum computer and observing that this would have been impossible without error mitigation. Then we propose the triamond lattice as an expedient approach to lattice gauge theories in three spatial dimensions, deriving the Hamiltonian and obtaining energy eigenvalues and eigenstates from a noiseless simulator for a three-dimensional unit cell.
I Introduction
@@
With quantum chromodynamics as a motivating long-range goal, the present work is focused on the simplest non-Abelian gauge theory. Being non-Abelian means that gauge fields carry the relevant charge directly and can therefore interact among themselves, like the gluons of quantum chromodynamics and in contrast to the photons of quantum electrodynamics. Specifically our work will consider SU(2) gauge theory in the absence of fermions, and we will truncate the gauge field to fit on a small number of qubits. Several research groups have already carried out exploratory studies of non-Abelian gauge theories on quantum computing hardware [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. There have also been many theoretical studies laying the groundwork for anticipated implementations on larger quantum computers [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. In the present work, we address two important issues: an effective way to mitigate errors and a practical way to extend lattices into three spatial dimensions.
@@
With successful mitigation in hand, we then turn our attention to lattice gauge theory in three spatial dimensions. While this can be done with a standard cubic lattice, it is important to remember that whenever more than three gauge links meet at a lattice site, the quantum numbers of the gauge links themselves are insufficient to fully define the state of the gauge field across the lattice [36, 40, 54, 58]. For example, consider four SU(2) gauge links meeting at a site and suppose each gauge link has quantum number $j=\frac{1}{2}$, where $j$ is notation familiar from quantized angular momentum though here it refers to the gauge degrees of freedom. One pair of these links could sum to $j_{\rm pair}=0$ or $1$. The other pair must have the same sum because Gauss’s law requires the total of all four to be $j_{\rm site}=0$. Therefore a full description of the state requires not just the $j$ values of individual links, but also the value of $j_{\rm pair}$ for some pair selected by the user to define that site.
One way to handle this issue is to assign a sufficient number of additional qubits at each lattice site, being careful to define a convention at each site, thereby completing the state’s definition [36, 40, 54, 58]. In the present work we propose an alternative based on a structure from crystallography [63, 64, 65, 66, 67, 68]. Having only three links at each site, our choice will avoid the need for any quantum numbers beyond the individual link values.
Consider a lattice where each site is touched by exactly three gauge links that have equal lengths, lie in a single plane, and are placed at equal angles around the site, i.e. $32\pi/$ 3. This forms a three-dimensional lattice that has a high degree of symmetry and needs no additional qubits beyond the links themselves. Because of its similarity to a diamond crystal of carbon atoms but with three links per site instead of four, it has been called the triamond lattice [66, 67]. It is also known as the Laves lattice [64] or the $K_{4}$ lattice [65, 68]. In this work, we provide an introduction to lattice gauge theory on the triamond framework. We derive the SU(2) Hamiltonian for a triamond lattice and demonstrate its use on a noiseless simulator.
II Results
II.1 Imaginary time evolution on square plaquettes
@@
Imaginary time evolution is used routinely in traditional lattice gauge theory on classical computers. Our motivation to study it now on a quantum computer is not to compete with successful classical methods, but rather to imagine that ground-state preparation will become the first step in a larger computation that truly does require quantum computing [69, 70].
Our computation uses a lattice comprising two side-by-side square plaquettes that share a single gauge link, thus forming a left path, a center path, and a right path as shown in Fig. 1. Each of the seven gauge links on this lattice is a superposition of SU(2) representations, and Gauss’s law requires the three on the left path to equal one another and similarly the three on the right path must equal one another. For our calculation we truncate the basis and retain only the two lowest states for each gauge link, $j=0$ and $j=\frac{1}{2}$.
Figure 1: The lattice used to demonstrate self-mitigation for quantum imaginary time evolution. The text refers to the 3 paths between sites A and B as the left path (three links), the center path (one link), and the right path (three links).
Any state of this system can be described by two qubits, one for the left path and one for the right path, because Gauss’s law then determines the center path uniquely. Such a tiny system is a valuable place to study error mitigation because a single time step requires only a simple circuit while more and more time steps will require more and more entangling gates. The real-time evolution of this system was studied previously as the first example of self-mitigation [17]. The Hamiltonian, in units of $2/g^{2}$, is
@@
The state tomography that determines the values of the coefficients $a_{jk}$ is presented in Sec. IV.1. With those values in hand, and with $r^{\prime}$ obtained from the calculation of Eq. (4), the state at a small value of $\tau$ can be computed from Eq. (3). In particular, we use
@@
where $CX_{jk}$ is a CNOT gate with qubit $j$ as control and qubit $k$ as target.
@@
Figure 2: The circuit used for one second-order Trotter step of imaginary time evolution for SU(2) gauge theory on a lattice comprising two square plaquettes.
Because the two-qubit CNOT gate is noisier than a single-qubit gate on quantum hardware, the gates in Fig. 2 have been ordered in a way that minimizes the number of CNOT gates. Also, the CNOT gate at the end of one Trotter step will cancel the CNOT gate at the beginning of the next step. As well, the very first CNOT gate in the circuit can be omitted because the initial control qubit is always off. Nevertheless, the open symbols in Fig. 3 show that only the first two or three time steps approach the true ground state energy and then subsequent time steps move further and further away. The quantum computation is overwhelmed by hardware noise without ever reaching the correct result.
Figure 3: Quantum imaginary time evolution, with and without self-mitigation, for SU(2) gauge theory with $j_{\rm max}=\frac{1}{2}$ on a lattice comprising two square plaquettes. The gauge coupling is $x=1$ and the time step is $\Delta\tau=0.1$ in units of $2/g^{2}$. Each quantum circuit uses 50 randomized compilings with $10^{4}$ shots per compiling. All data points were computed on ibm_lagos.
Self-mitigation is able to extract the true result from the computed data. The original implementation [17], which was inspired by the work of Urbanek et al. [72], has subsequently been used and extended in various ways [18, 19, 21, 56, 73, 74, 75, 76, 70, 77]. The basic idea is to run a mitigation circuit that is identical to the physics circuit except that $\tau\to-\tau$ in the second half of the circuit. If your circuit has an odd number of second-order Trotter steps, use $2\tau\to 0$ in the center gate (see, for example, Fig. 2) and insert a barrier to prevent CNOT cancellation.
@@
More precisely, randomized compiling (see Sec. IV.2) is used to convert CNOT errors into incoherent noise that is well described by the depolarizing noise model [72]. For our example of imaginary time evolution, this leads to a pair of ratios,
@@
The filled data points of Fig. 3 were obtained by computing each term of Eq. (1) from the ratios in Eqs. (9) and (10). The physics runs and mitigation runs (with 50 randomized compilings for each) and the readout error mitigation runs were all completed within a single qiskit job [78] to ensure they were experiencing the same hardware conditions. Whereas the unmitigated data points in Fig. 3 approach the pure noise limit of $\frac{21}{8}$, where only the first term in Eq. (1) makes a nonzero contribution, the self-mitigated data points in Fig. 3 remain consistent with the true ground state as $\tau$ increases. Error bars grow as the circuit gets longer, but meaningful results are still obtained with more than 400 CNOT gates. This success provides an incentive to move toward larger lattices, including novel three-dimensional approaches like the triamond lattice.
II.2 The SU(2) Hamiltonian for a triamond lattice
Figure 4: Two views of the unit cell for a triamond lattice. Each unit cell contains 8 sites and 12 gauge links.
Figure 5: The triamond lattice having 3 unit cells along each direction.
Figs. 4 and 5 show the locations of lattice sites and gauge links on a triamond lattice. See also our two-minute video presentation [79], which provides views of the triamond lattice from a camera moving around it.
@@ II.2 The SU(2) Hamiltonian for a triamond lattice
The angle between any two planes is $\arccos(\frac{1}{3})$, which is familiar because this same angle lies between any pair of lines joining the center of a cube to neighboring corners. Because the unit vector orthogonal to a plane is only unique up to its sign, the triamond lattice is chiral. Specifically, the mirror image of Fig. 5 is a distinct but equally valid triamond lattice.
@@
$H_{E}^{\rm cont}$ $=$ $\frac{1}{g^{2}}\int{\rm Tr}\left(E_{x}^{2}+E_{y}^{2}+E_{z}^{2}\right)d^{3}x\,,$ (14)
@@
where each trace is over the SU(2) indices of the electric field $\vec{E}$ or field strength tensor $F_{\mu\nu}$.
@@
$U(\vec{n},\hat{s})=e^{i\hat{s}\cdot\vec{A}(\vec{n})}\,,$ (16)
and the component of the gauge field at site $\vec{n}$ that points in direction $\hat{s}$ is an element of the SU(2) algebra,
@@
Our computations will be performed in the electric basis, where the magnetic terms are best expressed as a sum of all elementary plaquettes in this lattice and each plaquette is the trace of the ordered product of the 10 gauge links comprising that plaquette. The somewhat tedious derivation of this lattice Hamiltonian is outlined in Sec. IV.3. The result is
@@
$H_{E}$ $=$ $\frac{4\sqrt{2}a^{3}}{3g^{2}}\sum_{n={\rm links}}{\rm Tr}\left(E_{x}^{2}(n)+E_{y}^{2}(n)+E_{z}^{2}(n)\right)\,,~{}~{}~{}~{}~{}$ (19)
$H_{B}$ $=$ $-\frac{\sqrt{2}}{g^{2}a}\sum_{\vec{w}={\rm white}}\sum_{s=1}^{6}\mathcal{P}_{s}(\vec{w})\,,$ (20)
@@
Any eigenstate of $H_{E}$ is fully defined by providing the quantum numbers for every gauge link on the triamond lattice, $j_{1}$, $j_{2}$, $j_{3}$, …$j_{N}$. Our shorthand for an eigenstate is |$\left|\{j\}\right>$. The diagonal matrix elements of the Hamiltonian are
$\{j\}\right|H_{E}\left|\{j\}\right>=\frac{4\sqrt{2}}{3g^{2}a}\sum_{n=1}^{N}j_{n}(j_{n}+1)\,,$ (21)
@@
II.3 Computations on the triamond unit cell
@@
A simple way to begin exploring this theory is to use the variational method, which will find the lowest energy state among all states that couple to a selected ansatz. Fig. 6 shows three energy eigenstates obtained from three different ansatze, each having just one variational parameter. (We name the parameter $\theta$ in every case, but they are three different parameters.) Each of these examples is successful in finding a true eigenstate of the Hamiltonian. The 12-qubit quantum circuits that produced these results can be found in Sec. IV.4.
Figure 6: Three energy eigenvalues for one unit cell of the triamond lattice with periodic boundary conditions. The gauge coupling is $g=1$. Each data point was obtained by running qiskit code with $10^{4}$ shots on the ibm_qasm noiseless simulator. For comparison, red dashed lines show the true eigenvalues.
@@
Block diagonalizing the Hamiltonian is simple enough for such a small lattice, but phenomenological applications will require lattices with more than a single unit cell, and this is where the Hilbert space grows rapidly. A lattice with $N$ unit cells with have $12N$ gauge links. Truncating each link to $j\in\{0,\frac{1}{2}\}$ leads to $2^{12N}$ basis states, and the ground state is built from $2^{4N-2}$ of them. Larger lattices are where the variational method, imaginary time evolution, and other algorithms implemented with error mitigation on quantum computers have the potential to be of great value.
III Discussion

### v2

@@ Abstract
Lattice gauge theory should be able to address significant new scientific questions when implemented on quantum computers. In practice, error-mitigation techniques have already allowed encouraging progress on small lattices. In this work we focus on a truncated version of SU(2) gauge theory, which is a familiar non-Abelian step toward quantum chromodynamics. First, we demonstrate effective error mitigation for imaginary time evolution on a lattice having two square plaquettes, obtaining the ground state using an IBM quantum computer and observing that this would have been impossible without error mitigation. Then we propose the triamond lattice as an expedient approach to lattice gauge theories in three spatial dimensions and we derive the Hamiltonian. Finally, error-mitigated imaginary time evolution is applied to the three-dimensional triamond unit cell, and its ground state is obtained from an IBM quantum computer. Future work will want to relax the truncation on the gauge fields, and the triamond lattice is increasingly valuable for such studies.
@@
With quantum chromodynamics as a motivating long-range goal, the present work is focused on the simplest non-Abelian gauge theory. Being non-Abelian means that gauge fields carry the relevant charge directly and can therefore interact among themselves, like the gluons of quantum chromodynamics and in contrast to the photons of quantum electrodynamics. Specifically our work will consider SU(2) gauge theory in the absence of fermions, and we will truncate the gauge field to fit on a small number of qubits. Several research groups have already carried out exploratory studies of non-Abelian gauge theories on quantum computing hardware [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. There have also been many theoretical studies laying the groundwork for anticipated implementations on larger quantum computers [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 69, 68, 70, 71]. In the present work, we address two important issues: an effective way to mitigate errors and a practical way to extend lattices into three spatial dimensions.
@@
With successful mitigation in hand, we then turn our attention to lattice gauge theory in three spatial dimensions. While this can be done with a standard cubic lattice, where six gauge links touch each lattice site, there is an important issue to address. Specifically, for any lattice where more than three gauge links meet at a lattice site, the quantum numbers of the gauge links themselves are insufficient to fully define the state of the gauge field across the lattice [38, 42, 59, 64]. To understand why, consider a two-dimensional square lattice where four SU(2) gauge links meet at every site. At one particular site, suppose each of the four gauge links happen to have quantum number $j=\frac{1}{2}$, where $j$ is notation familiar from quantized angular momentum though here it refers to the gauge degrees of freedom. One pair of these links could sum to $j_{\rm pair}=0$ or $1$. The other pair must have the same sum because Gauss’s law requires the total of all four to be $j_{\rm site}=0$. Therefore a full description of the state requires not just the $j$ values of individual links, but also the value of $j_{\rm pair}$ for some pair selected by the user to define that site.
One way to handle this issue is to assign a sufficient number of additional qubits at each lattice site, being careful to define a convention at each site, thereby completing the state’s definition [38, 42, 59, 64]. The number of additional qubits needed to define $j_{\rm pair}$ grows logarithmically with the user’s choice for $j_{\rm max}$. Moreover, a three-dimensional lattice needs more than a single $j_{\rm pair}$ at each site. In the present work we propose an alternative based on a structure from crystallography [73, 74, 75, 76, 77, 78]. Having only three links at each site, our choice will avoid the need for any quantum numbers beyond the individual link values.
Consider a lattice where each site is touched by exactly three gauge links that have equal lengths, lie in a single plane, and are placed at equal angles around the site, i.e. $32\pi/$ 3. This forms a three-dimensional lattice that has a high degree of symmetry and needs no additional qubits beyond the links themselves. Because of its similarity to a diamond crystal of carbon atoms but with three links per site instead of more, it has been called the triamond lattice [76, 77]. It is also known as the Laves lattice [74] or the $K_{4}$ lattice [75, 78]. In this work, we provide an introduction to lattice gauge theory on the triamond framework. We derive the SU(2) Hamiltonian for a triamond lattice and demonstrate its use on a noiseless simulator. Then we employ our method for error-mitigated imaginary time evolution to create the ground state of the three-dimensional triamond unit cell on an IBM quantum computer [79].
Results
Imaginary time evolution on square plaquettes
@@ Results
Imaginary time evolution is used routinely in traditional lattice gauge theory on classical computers. Our motivation to study it now on a quantum computer is not to compete with successful classical methods, but rather to imagine that ground-state preparation will become the first step in a larger computation that truly does require quantum computing [80, 81, 82]. Imaginary time evolution has recently been applied to $Z_{2}$ gauge theory as well [83].
Our first computation uses a lattice comprising two side-by-side square plaquettes that share a single gauge link, thus forming a left path, a center path, and a right path as shown in Fig. 1. Each of the seven gauge links on this lattice is a superposition of SU(2) representations, and Gauss’s law requires the three on the left path to equal one another and similarly the three on the right path must equal one another. For our calculation we truncate the basis and retain only the two lowest states for each gauge link, $j=0$ and $j=\frac{1}{2}$.
Figure 1: A minimal lattice. This is the first lattice used to demonstrate self-mitigation for quantum imaginary time evolution. The text refers to the 3 paths between sites A and B as the left path (three links), the center path (one link), and the right path (three links).
Any state of this system can be described by two qubits, one for the left path and one for the right path, because Gauss’s law then determines the center path uniquely. Such a tiny system is a valuable place to study error mitigation because a single time step requires only a simple circuit while more and more time steps will require more and more entangling gates. The real-time evolution of this system was studied previously as the first example of self-mitigation [19]. The Hamiltonian, in units of $g^{2}/2$, is
@@
The state tomography that determines the values of the coefficients $a_{jk}$ is presented in Methods. With those values in hand, and with $r^{\prime}$ obtained from the calculation of Eq. (4), the state at a small value of $\tau$ can be computed from Eq. (3). In particular, we use
@@
where $CX_{jk}$ is a controlled not (CNOT) gate with qubit $j$ as control and qubit $k$ as target.
@@
Figure 2: One Trotter step. This circuit is used for one second-order Trotter step of imaginary time evolution for SU(2) gauge theory on a lattice comprising two square plaquettes. $\tau$ is the real-valued Euclidean time parameter. The coefficients $a_{jk}$ define time evolution as shown in Eq. (5).
Because the two-qubit CNOT gate is noisier than a single-qubit gate on quantum hardware, the gates in Fig. 2 have been ordered in a way that minimizes the number of CNOT gates. Also, the CNOT gate at the end of one Trotter step will cancel the CNOT gate at the beginning of the next step. As well, the very first CNOT gate in the circuit can be omitted because the initial control qubit is always off. Nevertheless, the error bars without symbols in Fig. 3 show that only the first two or three time steps approach the true ground state energy and then subsequent time steps move further and further away. The quantum computation is overwhelmed by hardware noise without ever reaching the correct result.
Figure 3: Quantum imaginary time evolution with and without self-mitigation. These results are for SU(2) gauge theory with $j_{\rm max}=\frac{1}{2}$ on a lattice comprising two square plaquettes. The gauge coupling is $x=1$ and the time step is $\Delta\tau=0.1$ in units of $2/g^{2}$. Each quantum circuit uses 50 randomized compilings with $10^{4}$ shots per compiling. All quantum computations were performed on ibm_lagos. Black data points use self-mitigation and blue data points do not. Error bars are 95% confidence intervals. The true ground state is shown as a red dashed line. The pure noise limit is shown as a blue dotted line.
Self-mitigation is able to extract the true result from the computed data. The original implementation [19], which was inspired by the work of Ref. [85], has subsequently been used and extended in various ways [20, 21, 23, 62, 86, 87, 88, 89, 81, 90]. The basic idea is to run a mitigation circuit that is identical to the physics circuit except that $\tau\to-\tau$ in the second half of the circuit. If your circuit has an odd number of second-order Trotter steps, use $2\tau\to 0$ in the center gate (see, for example, Fig. 2) and insert a barrier to prevent CNOT cancellation.
@@
More precisely, randomized compiling (see Methods) is used to convert CNOT errors into incoherent noise that is well described by the depolarizing noise model [85]. For our example of imaginary time evolution, this leads to a pair of ratios,
@@
The filled data points of Fig. 3 were obtained by computing each term of Eq. (1) from the ratios in Eqs. (9) and (10). The physics runs and mitigation runs (with 50 randomized compilings for each) and the readout error mitigation runs were all completed within a single qiskit job [91] to ensure they were experiencing the same hardware conditions. Whereas the unmitigated data in Fig. 3 approach the pure noise limit of $\frac{21}{8}$, where only the first term in Eq. (1) makes a nonzero contribution, the self-mitigated data points in Fig. 3 remain consistent with the true ground state as $\tau$ increases. Error bars grow as the circuit gets longer, but meaningful results are still obtained with more than 400 CNOT gates. This success provides an incentive to move toward larger lattices, including novel three-dimensional approaches like the triamond lattice.
The SU(2) Hamiltonian for a triamond lattice
Figure 4: The triamond lattice. Panel (a) shows two views of the unit cell, which always contains 8 sites and 12 gauge links. Panel (b) shows a lattice having 3 unit cells along each direction. In both panels, the colour of each link defines its direction according to Eqs. (11) and (12).
Fig. 4 shows the locations of lattice sites and gauge links on a triamond lattice. See also our two-minute video presentation (https://vimeo.com/904585432), which provides views of the triamond lattice from a camera moving around it.
@@
The angle between any two planes is $\arccos(\frac{1}{3})$, which is familiar because this same angle lies between any pair of lines joining the center of a cube to its corners. Because the unit vector orthogonal to a plane is only unique up to its sign, the triamond lattice is chiral. Specifically, the mirror image of Fig. 4 is a distinct but equally valid triamond lattice.
@@
$H_{E}^{\rm cont}$ $=$ $g^{2}\int{\rm Tr}\left(E_{x}^{2}+E_{y}^{2}+E_{z}^{2}\right)d^{3}x\,,$ (14)
@@
where each trace is over the SU(2) indices of the electric field $\vec{E}$ or field strength tensor $F_{\mu\nu}$. (Other normalization conventions are sometimes used for $\vec{E}$ and $F_{\mu\nu}$ [64].)
@@
$U(\vec{n},\hat{s})=e^{ia\hat{s}\cdot\vec{A}(\vec{n})}\,,$ (16)
where $a$ is the lattice spacing. The component of the gauge field at site $\vec{n}$ that points in direction $\hat{s}$ is an element of the SU(2) algebra,
@@
Our computations will be performed in the electric basis, where the magnetic terms are best expressed as a sum of all elementary plaquettes in this lattice and each plaquette is the trace of the ordered product of the 10 gauge links comprising that plaquette. The somewhat tedious derivation of this lattice Hamiltonian is outlined in Methods. The result is
@@
$H_{E}$ $=$ $\frac{8\sqrt{2}a^{3}g^{2}}{3}\!\!\!\sum_{n={\rm links}}{\rm Tr}\left(E_{x}^{2}(n)+E_{y}^{2}(n)+E_{z}^{2}(n)\right),~{}~{}~{}~{}~{}$ (19)
$H_{B}$ $=$ $-\frac{2\sqrt{2}}{g^{2}a}\sum_{\vec{w}={\rm white}}\sum_{s=1}^{6}\mathcal{P}_{s}(\vec{w})\,,$ (20)
@@
Any eigenstate of $H_{E}$ is fully defined by providing the quantum numbers for every gauge link on the triamond lattice, $j_{1}$, $j_{2}$, $j_{3}$, …, $j_{N}$. Our shorthand for an eigenstate is |$\left|\{j\}\right>$. The diagonal matrix elements of the Hamiltonian are
$\{j\}\right|H_{E}\left|\{j\}\right>=\frac{8\sqrt{2}g^{2}}{3a}\sum_{n=1}^{N}j_{n}(j_{n}+1)\,,$ (21)
@@
Computations on the triamond unit cell
@@
A simple way to begin exploring this theory is to use the variational method, which will find the lowest energy state among all states that couple to a selected ansatz. Fig. 5 shows three energy eigenstates obtained from three different ansatze, each having just one variational parameter. (We name the parameter $\theta$ in every case, but they are three different parameters.) Each of these examples is successful in finding a true eigenstate of the Hamiltonian. The 12-qubit quantum circuits that produced these results can be found in Methods.
@@
Figure 5: Three energy eigenvalues for one unit cell of the triamond lattice. Panels (a), (b) and (c) correspond to the third excited state, the first excited state, and the ground state respectively. The gauge coupling is $g=1$ and periodic boundary conditions are used in all directions. Each data point was obtained by running qiskit code with $10^{4}$ shots on the ibm_qasm noiseless simulator. Error bars are standard deviations. For comparison, red dashed lines show the true eigenvalues.
@@
Block diagonalizing the Hamiltonian is simple enough for such a small lattice, but phenomenological applications will require lattices with more than a single unit cell, and this is where the Hilbert space grows rapidly. A lattice with $N$ unit cells will have $12N$ gauge links. Truncating each link to $j\in\{0,\frac{1}{2}\}$ leads to $2^{12N}$ basis states, and the ground state is built from $2^{4N-2}$ of them. Larger lattices are where the variational method, imaginary time evolution, and other algorithms implemented with error mitigation on quantum computers have the potential to be of great value.
Triamond imaginary time evolution
When future computations are performed on triamond lattices, a typical first step will be to prepare the interacting vacuum. Imaginary time evolution is a reliable method for doing so, but it is only practical if error mitigation methods are able to account for the hardware errors. As a first confirmation of self-mitigation for triamond lattices, consider a single unit cell with periodic boundary conditions. The Hamiltonian matrix is Eq. (25), which can be written as
$\hat{H}_{\rm vac}$ $=$ $\frac{2\sqrt{2}g^{2}}{a}\bigg{(}6-2(\hat{Z}_{0}+\hat{Z}_{1})-2\hat{Z}_{1}\hat{Z}_{0}$ (26)
$-\frac{17}{64g^{4}}(\hat{X}_{0}+\hat{X}_{1}+\hat{X}_{1}\hat{X}_{0})$
$-\frac{15}{64g^{4}}(\hat{Z}_{1}\hat{X}_{0}+\hat{X}_{1}\hat{Z}_{0}-\hat{Y}_{1}\hat{Y}_{0})\bigg{)}\,.$
This is similar to the Hamiltonian for square plaquettes given in Eq. (1), but with different coefficients and additional terms. We adapted our code from the square plaquette study to handle the triamond case, obtaining the results displayed in Fig. 6. Perhaps surprisingly, the unmitigated results are moving away from the true value already at the very first step of imaginary time. Notice, however, that we are performing a particularly difficult computation: the pure noise limit is at a height of 6 (far beyond the top of the graph) while the true result is only slightly below zero. This means the true result contains a delicate cancellation between the large first term in Eq. (26) and the sum of all other terms.
Figure 6: Quantum imaginary time evolution on a triamond lattice. These results are for SU(2) gauge theory with $j_{\rm max}=\frac{1}{2}$ on a triamond unit cell with periodic boundaries in all three spatial directions. The gauge coupling is $g=1$ and the time step is $\Delta\tau=0.05$ in units of $a/(2\sqrt{2}g^{2})$. Each quantum circuit uses between 32 and 50 randomized compilings with 2000 shots per compiling. All quantum computations were performed on ibm_brisbane. Black data points use self-mitigation and blue data points do not. Error bars are 95% confidence intervals. The true ground state is shown as a red dashed line.
As expected, the unmitigated data points in Fig. 6 are approaching the pure noise limit as imaginary time increases. Fluctuations in that approach are observed and were expected because the computation at each time step was usually performed hours or days after the previous time step, and was run on whichever qubits were performing best at that time. In contrast, the self-mitigated results arrive at the correct value after two steps of imaginary time and remain in agreement for the duration of our computations. Even the fluctuations evident in the unmitigated results are handled nicely by self-mitigation, with no visible effect remaining in the self-mitigated data.
For $\tau<0.7$ in the units of Fig. 6, we used 50 randomized compilings. As $\tau$ increased beyond $0.7$, the number of randomized compilings was reduced linearly to just 32 randomized compilings at $\tau=1.1$. No sensitivity to this choice is observed in the results.
Table 1: The 16 options for randomized compiling. Each of the 16 rows in this table is equivalent to a single CNOT gate on error-free hardware. On noisy hardware, randomly selecting from these rows is valuable for error mitigation. See the discussion in Methods.
Applied before the CNOT Applied after the CNOT
control target control target
I I I I
I X I X
I Y Z Y
I Z Z Z
X I X X
X X X I
X Y Y Z
X Z Y Y
Y I Y X
Y X Y I
Y Y X Z
Y Z X Y
Z I Z I
Z X Z X
Z Y I Y
Z Z I Z
It is also worth noticing a significant difference between the two quantum computers we have used in this work. The square plaquette results in Fig. 3 ran on ibm_lagos where the CNOT gate is a native gate, but our triamond results in Fig. 6 ran on ibm_brisbane where CNOT gates are not native and are constructed from the echoed cross-resonance (ECR) gate. Self-mitigation has been tailored specifically to an ECR device by other authors [88] but our code is written in terms of CNOT gates (see Table 1), and it is encouraging to observe that a CNOT-based self-mitigation code performs well on the ECR device also.
Discussion